# Sensory Glove-Based Surgical Robot User Interface

### v1

@@ III-B Results
MLP 0.9852 0.9852 0.9852
L-GBM 0.9508 0.9505 0.9508
RF 0.9417 0.9413 0.9417
HGBT 0.9628 0.9627 0.9628
TABLE IV: Evaluation metrics for the trained models.
Although all models exhibited high accuracy, F1 scores, and recall, we selected the MLP model for implementation due to its high prediction speed (Table V) and high accuracy. The average frequency at which the MLP model can operate (determined from the ROS logs) is around 96Hz. From Fig. 2, it is noticeable that the Pinky and Ring gestures are associated with more errors. This observation can be attributed to the similarity between these two gestures.
(a) MLP
(b) L-GBM
(c) Random Forest
(d) HGBT
Figure 2: Confusion matrices for gesture recognition models.
@@
MLP 0.75 0.14
L-GBM 2.09 3.06
RF 0.92 0.16
HGBT 7.81 4.72
TABLE V: Prediction time for each model (1000 trials).
In our application, gesture recognition needs to be performed continuously and in real-time. To ensure the stability of recognition – for example, even a single mistake in the clutch could lead to problematic input sent to the robot – we use a sliding window of length 7. The predictions of the MLP model over the last 7 time steps are stored in a 7x5 matrix, with each row corresponding to a time sample and each column representing a gesture prediction (0 or 1). The matrix is initialized with zeros. During each iteration, a new row is added with the most probable gesture set to one; the oldest prediction is deleted from the matrix. Finally, in each iteration, the system selects the gesture with the highest column sum. This induces a negligible delay but greatly increases the stability of the prediction.
IV Preliminary Evaluation
To evaluate the system, the members of the team with surgical training were asked to use the system to complete one of the standard da Vinci training exercises [26]. The original exercise was modified to suit our system’s single-arm setup and to simplify the task, thereby reducing any uncertainties in the preliminary test. The modified task required participants to control the PSM with their left hand to lift a peg from one pin on the pegboard and transfer it to another pin. One trial consists of repeating this task twice with two different pegs on the pegboard.
IV-A Experiment Setup
The experimental setup is shown in Fig. 1. The user’s hand starts in a position similar to that shown in the figure, away from the robot. The endoscope was positioned to allow a view of the pegs on the board and the tip of the instrument. God Vision glasses were connected to the vision tower, and the user wore the glasses to access the stereo endoscopic view. The evaluation was carried out by 5 surgeons and one fellow, all members of the research team. All have prior experience with the da Vinci surgical console to perform surgical tasks. Participants underwent a brief training session, lasting approximately 2 to 10 minutes, primarily to familiarize themselves with the operation of the clutch mechanism. Each participant performed the exercise 5 times. To implement our system, we needed to update the dVRK. This caused compatibility issues with the existing dVRK console so we were not able to conduct the exercise through the standard da Vinci console.
@@ IV-B Control Results
(c) Jaw Movements
Figure 3: A sample recording of motion between the glove and the PSM. The segments highlighted in purple indicate periods when the clutch mechanism was engaged. (a) displays the translation of the glove and PSM while (b) shows the orientation of the glove and PSM respectively. (c) compares the distance of the thumb and index finger on the glove with the jaw angle of the PSM.
Figure 4: Translational and orientational distance between the glove and the PSM: the left figure depicts the Euclidean distance, while the right one details the orientation difference, quantified with rotation matrices (in the hand workspace).
Figure 5: Both plots detail the x-axis translation of the glove and the PSM tip in Fig. 2(a). The initial plot illustrates a delay in the robot’s execution of the glove’s movements. To address this, the delay between the signals was measured based on the peaks, and the aligned signals were shown on the bottom.
A sample trajectory of the da Vinci arm controlled with our proposed glove, performed by one of the surgeons, is shown in Fig. 3. A comparison of the kinematics of the glove and the instrument tip is shown in Figs. 2(a) and 2(b), respectively. The motion of the PSM exhibits a stair-case movement due to the implementation of position control of the arm in dVRK (green box in Fig. 5), where the robot temporarily stops once it reaches the goal position. The shaded regions show the moments when the clutch (Fist gesture) was activated. Our clutch mechanism (Section II-G) prevents the PSM from moving while activated and resumes the movement smoothly when released. The motion of the index and thumb of the glove and the instrument jaw is depicted in Fig. 2(c). Currently, the fully closed configuration of the FBF jaw is represented by a negative value in the dVRK’s system (as the $0^{\circ}$ angle corresponds to a parallel configuration of the jaw where it is not completely closed); the slight difference for extreme values between the glove and the PSM is due to the scaling between the glove and the jaw.
To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (the inverse of Eq. (1)) of the movement of the instrument tip (Fig. 5). The figure clearly shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics. To accurately estimate the error between the input and output signals and eliminate the effect of the delay, we identified the delay by the peak positions within each signal. Subsequently, the PSM signal is shifted according to the calculated delay to align with the glove’s signal (Fig. 5). The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [27]. The errors derived from Figs. 2(a) and 2(b) are shown in Fig. 4. The duration and error metrics for each participant’s performance are depicted in Table VI. The average delay in all trials and all participants was 0.223s.
User Average Duration (s)
@@
Error (rad)
Average Std. dev. Average Std. dev.
A 60.345 0.005 0.004 0.024 0.024
B 45.304 0.004 0.004 0.016 0.018
C 57.922 0.005 0.004 0.040 0.056
D 53.298 0.004 0.004 0.021 0.026
E 52.506 0.003 0.003 0.015 0.015
F 37.782 0.003 0.003 0.021 0.021
TABLE VI: the first column displays the average runtime for the five trials. The subsequent columns detail the overall mean and standard deviation of the errors, calculated based on the glove’s workspace, performed by each participant (complete version available on [28]).
@@ IV-C Survey Results
In the follow-up survey, each participant was asked five questions: 1) How comfortable was the glove controller? 2) How responsive was the robot when controlled with the glove controller? 3) How intuitive was controlling the jaw with your fingers? 4) How intuitive was it to control the arm with the glove? 5) Was the vibrotactile feedback for the clutch effective? We use a 7-point Likert scale for the response, with 7 being the most positive response [29]. The average responses for the 6 participants are summarized in Table VII. The results of Questions 3 to 5 show that the new control architecture is intuitive. This is further evidenced by the small amount of practice required to complete the exercise. Questions 1 and 2 produced satisfactory results; all participants responded that the system was ’highly responsive.’ However, the results of Question 1 indicate that the overall ergonomics of the system can be improved, as most users had to lift their elbow to achieve a correct angle with their hand. This issue also highlights two important aspects: first, the user’s experience with the device may significantly impact their comfort, as they can readjust themselves by clutching to find a better position. Second, an ergonomic seat can greatly reduce the user’s effort.
Question # Q1 Q2 Q3 Q4 Q5
Avg. Score 5.50 6.00 6.50 6.50 6.33
TABLE VII: Average scores in the Likert scale from the users.
@@ V Conclusion
This paper presents a novel user interface for the da Vinci surgical robot arm. The system integrates the Manus Meta Prime 3 Haptic XR glove, the HTC Vive Tracker, God Vision wireless smart glasses, and controls one arm of the dVRK surgical robot platform. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. A custom robot calibration was implemented to address the inherent inaccuracies of the dVRK system. A combination of scaling and gesture recognition was used to achieve precise and intuitive control of the robot’s movements while implementing an upgraded version of the clutch mechanism. In particular, we introduced clutching of the instrument orientation, a functionality not available in the da Vinci system. The vibrotactile elements of the glove are used to provide feedback to the user when gesture commands are invoked.
@@
The proposed interface offers several potential advantages. Surgeons can directly control robotic arms by moving their hands, without the need for additional physical devices. This facilitates a more immersive surgeon-robot interaction. However, it should be noted that the hardware used in our setup is not the latest available on the market; we used the first version of HTC Vive Trackers and Manus Meta Prime 3 XR, both of which have lower precision compared to their upgraded versions (HTC Vive Trackers v2 and Quantum XR Metaglove).
@@
The system was tested by members of the team with surgical training, including 5 surgeons and 1 fellow. They were asked to perform a surgical training task similar to those used in training with the da Vinci devices. The evaluation shows that the system is responsive and precise. Surgeons were able to effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. After the exercise, each participant filled out a questionnaire to provide feedback on their user experience. The responses show a high level of user satisfaction.
@@
Existing surgical consoles are bulky and take up valuable space. They limit the ability to optimize the operating room setup for throughput. They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions. In contrast, our proposed system is inexpensive as it uses off-the-shelf components, and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
@@
Future work includes implementing velocity control to enhance user experience, upgrading hardware to improve tracking precision, enabling full control for the right hand, and expanding the repertoire of user gestures to increase the range of available actions.

### v2

@@ III-B Training Results
FNN 0.9857 0.9764 0.9784
L-GBM 0.9875 0.9873 0.9875
RF 0.9830 0.9826 0.9830
HGBT 0.9792 0.9789 0.9792
TABLE III: Evaluation metrics of the trained models.
Table III shows the evaluation results on the test set. Based on the metrics, all models had analogous performance in estimating the hand gestures. Hence, we conducted an additional live test in which the models predicted the gestures in real time for 1000 iterations each. As shown in Table IV, L-GBM was the fastest among the models with an average delay of 1.63 ms. The speed is mainly due to its efficient way of growing trees and making their structure as compact as possible. The hyperparameters that were used for training the current L-GBM model are as follows: learning_rate=0.1, max_depth=15, min_child_samples=10, n_estimators=100, and num_leaves=15.
When using the classifiers in real-time applications, the classifier’s output is often post-processed using a majority voting technique [32]. While this introduces fixed time delays depending on voting buffer size, it enhances the accuracy and reduces short-term false transitions. For our application, we set the voting buffer size to 20. Thus, the predictions of the L-GBM model over the last 20 iterations are stored in a $20\times 3$ matrix, where each row contains the classifier estimations for each gesture. The gesture (column) with the highest sum is selected at the end.
@@
FNN 2.70 0.95
L-GBM 1.63 1.59
RF 3.90 0.62
HGBT 4.54 1.40
TABLE IV: Real-time prediction time for each model (1000 trials).
IV Experimental Results
To evaluate our system, 5 surgeons and 2 medical doctors with experience in using the da Vinci surgical console for surgical tasks were asked to use the system to complete one of the standard da Vinci training exercises [33] with both the console and the sensory gloves. The original exercise was modified to suit our system’s single-arm setup. The modified task required participants to control the PSM with their left hand to lift a peg from one pin on the pegboard and transfer it to another pin. One trial repeats this task twice with two different pegs on the pegboard.
IV-A Experimental Setup
The experimental setup is shown in Fig. 1. The user’s hand starts in a position similar to that shown in the figure, away from the robot. The endoscope was positioned to allow a view of the pegs on the board and the tip of the instrument. The SCOPEYE smart glasses were connected to the vision tower, providing a stereo endoscopic view to the user. Participants underwent a brief training session, lasting approximately 2 to 10 minutes, primarily to familiarize themselves with the glove and the operation of the clutch mechanism. Each participant performed the exercise 5 times, both with the glove system and with the dVRK console.
@@ IV-B Control Results
(c) Jaw Movement
Figure 2: A sample trajectory of the glove and the PSM: (a) translation; (b) orientation; and (c) the distance of the thumb and index fingers on the glove and the jaw angle of the PSM. The gray and purple regions correspond to the periods when Ring and Clutch gestures, respectively, are active.
A sample trajectory of the da Vinci arm controlled by one of the surgeons with our proposed system is shown in Fig. 2. A comparison of the kinematics of the glove and the instrument tip is shown in Figs. 2(a) and 2(b), respectively. The motion of the PSM exhibits a staircase movement due to the implementation of position control of the arm in dVRK (green box in Fig. 3), where the robot stops once it reaches the reference position.
In Fig. 2, the shaded region in gray (3 seconds) corresponds to the initialization of the system. This occurs after the Ring gesture has been activated for two seconds. During initialization, the arm tries to align with the pose of the glove. The shaded region in purple corresponds to the period during which the Clutch was activated. The Clutch prevents the PSM from moving; the motion is resumed when the Clutch is released. The motion of the index and thumb of the glove and the instrument jaw is depicted in Fig. 2(c).
Figure 3: The x-axis translation of the glove and the PSM tip: raw data (top) and data after the delay has been eliminated (bottom).
Controller
@@
Error (rads)
Delay (s)
Mean Std. Dev. Mean Std. Dev.
Glove 0.005 0.010 0.096 0.146 0.205
Console 0.002 0.001 0.038 0.066 0.001
TABLE V: The average translational and rotational errors between each controller and the PSM across all participants. The last column shows the average delay between the controller and the PSM tip signals.
User Glove Console
Avg. Time
Avg. Travel
Distance (m)
Avg. Time
Avg. Travel
Distance (m)
A 42.7 0.503 51.9 0.585
B 40.3 0.541 35.8 0.479
C 33.5 0.442 29.2 0.375
D 39.3 0.606 18.5 0.298
E 50.2 0.726 30.8 0.365
F 38. 0.479 27.5 0.302
G 34.3 0.503 27.1 0.382
Total 39.8 0.543 31.6 0.398
TABLE VI: Performance evaluation. The first column displays the average runtime, and the subsequent column shows the average traveled Euclidean distance of the instrument tip manipulated by each controller of the five trials from each participant.
To compare the input motion of the glove and the output motion of the instrument tip, we first reversed the scaling (Eq. (1)) of the movement of the instrument tip (Fig. 3). The figure shows a delay between the two trajectories, mainly due to the computation time for the inverse kinematics. To accurately estimate the error between the input and output signals, we identified the delay by the peak positions within each signal and eliminated the effect of the delay by shifting the PSM signal to align with the glove’s signal (Fig. 3). The difference in translation was quantified using the Euclidean distance, while the discrepancy in rotation was determined using the standard bi-invariant metric on the group of rotations [34]. The errors for each controller in all the trials across the participants are listed in Table V. Although the magnitude of the errors is larger for our system, the task completion times in the two setups were comparable (Table VI), suggesting that the errors were acceptable. These larger errors might also explain the slightly higher average travel distances for our system.
The main reason for significantly higher std. deviation for our system is mainly the consequence of safety mechanisms in the dVRK console: when the difference between the PSM position and the console position is larger than a threshold (as when the PSM is held by a peg, for example), the system shuts down. No such mechanism is employed in our system, leading to the presence of large errors.
The average delay of our system is $0.2$s, significantly larger than for the dVRK console. However, [35] shows that such delays are routinely tolerated by humans.
We evaluated the average performance of the participants with each controller (Table VI). Performance is evaluated by the average duration and travel distance of the instrument tip, indicating how efficiently they managed to perform the task. Considering that this is a prototype system, the performance of controlling the robot with the glove compared favorably with that of the console.
@@ IV-C Survey Results
Participants were asked to compare the usability and task load of our proposed system with the conventional console. Usability was assessed with the System Usability Scale (SUS) [17], and task load was measured using NASA Task Load Index (TLX) [18] questionnaires. The overall results are shown in Table VII.
The average SUS score for the glove was 67.5 compared to the console’s 81.4 (13.9 points of difference). These results align well with the quantitative assessment of the performance in Table VI: users who achieved similar performance between the glove and the console responded with a similar SUS assessment for the two. Considering that users have significantly more training on the dVRK console and that our system is just a prototype, we judge the reported SUS score as overwhelmingly positive.
When considering the workload, the average TLX score of the glove was higher than that of the console (40.6 vs. 26.1). We hypothesize that this is mainly due to the lack of ergonomic support in our system, while the dVRK console is equipped with a proper armrest and headrest. In addition, as in the case of the SUS scores, users are more familiar with the da Vinci console setup, significantly reducing the effort needed to adapt to the system. Thus, we anticipate that these scores would improve if users spent more time with our system.
@@ V Conclusion
This paper presents a novel surgeon interface for the da Vinci surgical robot. The system integrates the Manus Meta Prime 3 Haptic XR glove, the HTC Vive Tracker, and SCOPEYE wireless smart glasses and controls one arm of the dVRK surgical robot platform. In addition to moving the arm, the surgeon can use fingers to control the end-effector of the surgical instrument. A combination of scaling and gesture recognition was used to achieve precise and intuitive control of the robot’s movements while implementing an enhanced version of the clutch mechanism. In particular, we introduce clutching of the instrument orientation, a functionality unavailable in the da Vinci system. The vibrotactile feedback of the glove is used to alert the user when gesture commands are invoked.
@@
The proposed interface offers several potential advantages. Surgeons can directly control robotic arms by moving their hands without the need for additional physical devices. This facilitates a more immersive surgeon-robot interaction. However, it should be noted that the hardware used in our setup is not the latest available on the market; we used the first version of HTC Vive Trackers and Manus Meta Prime 3 XR, both of which have lower precision compared to the next generation devices (HTC Vive Trackers v2 and Quantum XR Metaglove).
Controller
Glove Console
User SUS TLX SUS TLX
A 92.5 27 92.5 35
B 70 45 85 27
C 67.5 45 60 65.8
D 80 10.8 80 5
E 22.5 60.8 92.5 11.6
F 82.5 48 92.5 12.5
G 57.5 47 67.5 25
Total Average 67.5 40.6 81.4 26.1
TABLE VII: System Usability Scale (SUS) and NASA Task Load Index (TLX) questionnaire results, answered by surgeons (green letters) and medical doctors (red letters). In the SUS, a score closer to 100 indicates better usability, and in the TLX, a score closer to 100 indicates a higher workload.
@@
The system was tested by users who were experienced with the da Vinci system, including five surgeons and two medical doctors. They were asked to perform a surgical training task similar to those used in training with the da Vinci devices. The evaluation shows that the system is responsive and precise. Surgeons could effectively complete training exercises with minimal practice with the new interface, suggesting that the interface is highly intuitive. After the exercise, each participant completed two questionnaires to assess the usability and workload of the proposed system. Users rated their experience favorably; however, a higher workload suggests the need for a more ergonomic setup.
@@
The existing surgical consoles are bulky and take up valuable space in the operating room. They limit the ability to optimize the operating room setup for throughput. They also pose challenges to the communication of the surgical team, leading to potential workflow disruptions. In contrast, our proposed system is inexpensive, as it uses off-the-shelf components, and it dramatically reduces space requirements. It demonstrates how modalities such as vibrotactile feedback can efficiently provide additional information to the surgeon without cluttering the visual field. Finally, the proposed architecture allows for rapid prototyping and opens opportunities for further innovations in the design of surgical robot interfaces.
@@
Future work includes implementing velocity control to enhance the user experience, enabling full control of both surgical arms,upgrading hardware to improve tracking precision, and expanding the repertoire of user gestures to increase the range of available actions.